OpenStudy (anonymous):

a particle execute shm along a straight line . the amplitude of oscillation is 2cm. when displacement of particle from the mean position is 1cm the magnitude of its acceleration is equal to magnitude of its velocity . the time period of oscillation is?

OpenStudy (cybershadow):

As we know that A = Amplitude = 2 cm x =displacement = 1 cm a = acceleration =-w^2x v =velocity = w (A^2-x^2)^1/2 as Magnitude of a = Magnitude of b i.e w^2x = w (A^2-x^2)^1/2 implies wx = (A^2-x^2)^1/2 also 2 = 2pi/T therefore 2pi/T = (A^2-x^2)1/2 T = 2pi/(A^2-x^2)1/2 T = 2pi/(4-1)^1/2 T = 2pi/sqrt(3) T = 1.73 seconds Check if this answer is correct , keep me posted :)